The output of the step down transformer is 48V when connected to the 12W bulb. What is the value of the peak current ?
 

Grade:12

1 Answers

Arun
5 years ago
Dear Akshay
 
Output/secondary voltage Vs = 48 Volt
Power associated with secondary Ps = 12 watt
Is = Ps/Vs = 12/48 = 0.25 amp
Peak current I0 = Is * sqrt (2) = 0.25 * sqrt(2)
 
Regards
Arun (askIITians forum expert)